OpenStudy (anonymous):

help me in this problem: f(x) denotes a function defined by sum of the digits of the number x it is given, f(a)<5 f(b)<5 f(c)<5 find the possible set of a,b,c for which f(a+b+c)>50

OpenStudy (amistre64):

what our domain of definition? reals, ints, wholes ....

OpenStudy (anonymous):

natural numbers..

OpenStudy (amistre64):

the sum of the digits, does that mean: f(a) = 1+2+3+...+a < 5 ?

OpenStudy (amistre64):

or say a=23; f(a) = 2+3 < 5

OpenStudy (anonymous):

the second one,@ amistre64

OpenStudy (anonymous):

natural numbers are positive integers, @ 2bornot2b

OpenStudy (amistre64):

n = {1,2,3,...}

OpenStudy (anonymous):

yeah, grade 11

OpenStudy (anonymous):

i see..

OpenStudy (amistre64):

our options for less than 100: a,b,c = {1,2,3,4,10,11,12,13,20,21,22,30,31,40} this is not all of the ns of course, just the first few that I see that sum < 5

OpenStudy (anonymous):

X + 4<-9 = 13 but how do you solve it?

OpenStudy (anonymous):

the main problem is f(a+b+c) >50....

OpenStudy (anonymous):

@2bor.. at least my school teaches that in class 11 :)

OpenStudy (anonymous):

U r in which grade?? and one thing, try out my problem too..

OpenStudy (amistre64):

by adding in 1 for 100; we get more: {110,111,112,120,121,130} same for 2 {210,211,220} {310} just getting a feel for it ; does f(a+b+c) mean 13+40+1 as: 1+3+4+0+1 ? or does it mean: 13+40+1 > 50 or: 13+40+1 = 54 = 5+4 > 50 ??

OpenStudy (anonymous):

@amistre, the second process, again :)